OpenStudy (anonymous):

Good morning all my math wizards:) Can someone tell me if I did this right in finding the product? (xy-5)^2 =(x^2)(y^2)-5^2=xy^4-5^2?

OpenStudy (hoblos):

the formula says: (a-b)^2 = a^2 - 2ab + b^2 so (xy-5)^2 = (xy)^2 - 10xy +25

OpenStudy (anonymous):

Oh man, I was way off. Ok, let me try it. I mean I am sure you are right, just have to do it to make sure I understand what I am doing;) ty sooooooooo much!!! Denise;)

OpenStudy (anonymous):

Where does the 10xy come from Hoblos?

OpenStudy (hoblos):

look at the formula i gave you.... 2ab= 2(xy)(5)=10xy

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I am sorry, I don't think I'm awake enough to be doing this but I will keep looking at the formula;) ty again, Denise:)

OpenStudy (hoblos):

my pleasure XD